OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sine of pi divided by two times cosine of pi divided by seven plus cosine of pi divided by two times sine of pi divided by seven.

OpenStudy (anonymous):

\[\sin \frac{ \pi }{ 2 }\cos \frac{ \pi }{ 7} + \cos \frac{ \pi }{ 2 }\sin \frac{ \pi }{ 7 }\]

OpenStudy (jdoe0001):

http://www.sosmath.com/trig/addform/img1.gif <--- notice the 3rd identity \(\bf \sin \frac{ \pi }{ 2 }\cos \frac{ \pi }{ 7} + \cos \frac{ \pi }{ 2 }\sin \frac{ \pi }{ 7 }\implies sin\left(\frac{ \pi }{ 2 }+\frac{ \pi }{ 7}\right)\)

OpenStudy (anonymous):

@jdoe0001 Thank you so much! that makes alot of sense to me now!